OpenStudy (please.help.me):

cos^3 x = cos x Find all solutions of the equation in the interval [0, 2π). (Enter your answers as a comma-separated list.

zepdrix (zepdrix):

\[\large\rm \cos^3x=\cos x\]Subtract cosx from each side,\[\large\rm \cos^3x-\cos x=0\]Next, factor a cosx out of each term.

zepdrix (zepdrix):

Can you tell me what that would look like?

OpenStudy (please.help.me):

cos x (cos x +1) (cos x -1)

zepdrix (zepdrix):

So you applied your difference of squares after factoring the cosx out? Ahh good good.

zepdrix (zepdrix):

\[\large\rm \cos x(\cos x+1)(\cos x-1)=0\]Next apply your Zero-Factor Property,\[\large\rm \cos x=0\]\[\large\rm (\cos x+1)=0\]\[\large\rm (\cos x-1)=0\]And solve for x in each of these cases.

OpenStudy (please.help.me):

Is the answer pi and 2pi?

zepdrix (zepdrix):

Taking this third part, \[\large\rm \cos x-1=0\]adding 1,\[\large\rm \cos x=1\]Cosine gives us 1 at an angle of 0. So x=0 is a solution. Which means, yes, 2pi would also be a solution (because 0 and 2pi are co-terminal). But notice that 2pi is not in our interval. When we have a rounded bracket on the interval, it means we `exclude` that particular end value.

zepdrix (zepdrix):

\[\large\rm \cos x+1=0\]subtracting 1,\[\large\rm \cos x=-1\]Cosine is negative 1 when we're on the left side of the x-axis, so yes pi.

OpenStudy (please.help.me):

So the answer is just pi?

zepdrix (zepdrix):

There is another solution or two coming from this though,\[\large\rm \cos x=0\]

zepdrix (zepdrix):

No... \(\large\rm (\cos x-1)=0\) gave us a solution. Read again what I posted if you're not sure. \(\large\rm (\cos x+1)=0\) gave us another solution. And \(\large\rm \cos x=0\) will give us two more solutions.

zepdrix (zepdrix):

So we should end up with 4 in total.

OpenStudy (please.help.me):

3pi/2 ?

OpenStudy (please.help.me):

Is that one of the solutions?

zepdrix (zepdrix):

Maybe, where did you get that from? :d

OpenStudy (please.help.me):

I am looking at the unit circle and cos is 3pi/2 at (0,-1)

OpenStudy (please.help.me):

So I know that pi and 3pi/2 are solutions. But I am confused how to get the other two, since you said there are 4 in all

zepdrix (zepdrix):

cosine is your x-coordinate. So at 3pi/2 you're getting the 0 from that coordinate pair. Ok good, so that's one of our solutions!

zepdrix (zepdrix):

So we want anywhere that cosine(the x-coordinate) is giving us zero. That would include up here at this point, ya?